# OpenWorldSAM: Extending SAM2 for Universal Image Segmentation with Language Prompts

### v3

@@ 4.3 Referring Expression Segmentation Performance Analysis
Table 3: Referring segmentation performance (cIoU) comparison on RefCOCOg benchmark validation set between our proposed OpenWorldSAM and prior SOTA methods. We abbreviate the datasets: C (COCO), RC (RefCOCO/+), RCg (RefCOCOg), PL (PACO-LVIS), O365 (Objects365), V (Video segmentation datasets), OID (OpenImages Detection), VG (Visual Genome), ADE (ADE20K), $P$ (PASCAL-Part), PC (PASCAL-VOC). We compare model trainable parameters, model capabilities (OV seg (open-vocabulary segmentation) and Inter Seg (interactive segmentation)), and training data required. “*” denotes an estimate of the trainable parameters, since these models use LoRA [hu2022lora] with rank-8/16 adapters for finetuning.
@@
X-Decoder (L) [zou2023xdecoder] CLIP-B [radford2021clip] (63M) 38 (M) C, RCg, Cap4M 64.6
SEEM (L) [zou2023seem] CLIP-B [radford2021clip] (63M) 39 (M) C, RC, RCg, PL 65.6
PolyFormer (L) [liu2023polyformer] BERT-B [devlin2019bert] (104M) 450 (M) RC, RCg 71.2
UNINEXT (H) [yan2023universal] BERT-B [devlin2019bert] (104M) 673 (M) C, RC, O365, V 74.4
APE-B (L) [shen2024ape] CLIP-L [radford2021clip] (123M) 42 (M) C, PC, O365, OID, VG, RC, RCg 63.5
PixelLM [ren2024pixellm] LLaMA2 [touvron2023llama2] (13B) 29 (M) C, RC, ADE, PL, MUSE 69.3
LISA [lai2024lisa] Vicuna [vicuna2023] (7B) 32 (M) C, RC, ADE, PL, PP 66.4
GLaMM [rasheed2024glamm] Vicuna [vicuna2023] (7B) 40 (M) RC, GranD 74.2
u-LLaVA [xu2024ullava] Vicuna [vicuna2023] (7B) 44 (M) C, RC, ADE, PL, PC 71.6
u-LLaVA [xu2024ullava] Vicuna [vicuna2023] (7B) 7 (B) C, RC, ADE, PL, PC 74.8
Sa2VA [yuan2025sa2va] InternVL2 [chen2024far] (1B) 22 (M) RC, RCg, V, GranD 72.3
Sa2VA [yuan2025sa2va] InternVL2 [chen2024far] (4B) 35 (M) RC, RCg, V, GranD 74.1
EVF-SAM [zhang2024evfsam] BEIT-3-L [wang2023beit] (0.7B) 674 (M) RC 77.0
OpenWorldSAM BEIT-3-L [wang2023beit] (0.7B) 4.5 (M) C, RCg 74.0
@@
Table 4: Ablation on the VLM choice, e.g., CLIP [radford2021clip] model from OpenAI. ‘✔’ denotes modality used during training, and ‘✘’ unused. Only the adapter modules are trainable, and the VLMs are kept frozen. Late fusion means we concatenate text/image features from the last layers of CLIP’s text/image encoder. Early fusion means BEiT-3 processes both modalities in all 24 Transformer layers.
@@ 4.4 Ablation Studies
Multi-modal encoder analysis. In Table 4, we compares performances using different VLM encoders and fusion methods (early fusion vs. late fusion). BEiT-3’s early cross-modal fusion (joint text-image processing across all layers) outperforms CLIP’s late fusion (last-layer concatenation) by +33.9 mIoU, +21.2 PQ, and +13.3 AP on ADE-150, demonstrating that deep semantic integration is critical for aligning language concepts with visual regions, echoing findings by EVF-SAM [zhang2024evfsam].
@@ A.1 Pre-training
We implement our model in PyTorch, building on the Detectron2 [wu2019detectron2] framework. We initialize the base models with the public weights of SAM2-Hiera-Large [^1] https://github.com/facebookresearch/sam2 and BEIT-3-Large [^2] https://huggingface.co/YxZhang/evf-sam2-multitask. The model is pre‑trained for 25 epochs on the COCO‑2017 training split (104K images) [lin2014coco]. We use the panoptic annotations, which provide pixel‑accurate masks and category labels for all 132 thing and stuff classes. Training is conducted on a single NVIDIA A100 (80 GB) GPU with a batch size of 8. Optimization employs AdamW (learning rate 1e-4). A step decay scheduler drops the learning rate by a factor of 0.1 at 89% and 96% of the total iterations. Compared with recent generalist models, our recipe is markedly more data‑efficient (see Table 6).
@@
X-Decoder [zou2023xdecoder] COCO, RefC Cap4M 32, 1024 200M (50 Ep $\times$ 4M Img)
OpenSeeD [zhang2023openseed] COCO, O365 32, 64 48M (30 Ep $\times$ 1.8M Img)
APE (B) [shen2024ape] COCO, LVIS, O365, OID, VG, RefC 16 17.28M (16 Bs $\times$ 1.08M Iter)
@@ A.2 Zero-Shot Evaluation
We evaluate semantic, instance, and panoptic segmentation in a zero‑shot setting. For instance segmentation and panoptic segmentation, we apply confidence-score filtering to remove masks with scores below 0.7, followed by non‑maximum suppression (NMS) with IoU threshold 0.5 to remove duplicate detections and retain distinct object instances. The confidence scores, originally termed “estimated IoU scores” in SAM [kirillov2023sam, ravi2024sam2], are direct outputs from SAM2’s mask decoder. These scores were optimized during SAM2’s pre-training to select high-quality (i.e., confident) mask outputs.
@@ Appendix C Additional Zero‑Shot Qualitative Results
Figure 8 showcases multiple challenging indoor scenes drawn from ADE20K-150/857 [zhou2017ade20k], and PASCAL Context-459 [mottaghi2014pascal-context]. In each sub-panel, we compare example outputs of OpenWorldSAM with comparisons to X-Decoder under both global-matching and oracle-prompts evaluation protocols.
@@
Figure 8: Qualitative comparisons between X-Decoder [zou2023xdecoder] and OpenWorldSAM on ADE20K-150, ADE20K-857, and PASCAL Context-459.
@@ Appendix D Limitations - Outdoor Generalization
Despite strong results on indoor and everyday photographs, OpenWorldSAM under-performs on driving datasets such as Cityscapes [cordts2015cityscapes] and BDD10K [yu2018bdd100k] (Table 8). Fine-tuning on Cityscapes narrows the gap, yet performance still trails methods explicitly exposed to multi-domain data. Understanding the source of this shortfall is essential for future extensions.
@@
X-Decoder (L) [zou2023xdecoder] zero-shot 52.0 24.9 38.1 47.2 17.8
@@
    Domain shift. COCO images are mostly handheld and indoor, whereas Cityscapes/BDD10K contain forward-looking dash-cam frames with motion blur, glare and night scenes. X-Decoder was co-trained on web-scale image-text pairs that include many outdoor photos, so its visual encoder has wider coverage. Large-scale multi-domain training is known to mitigate domain shift [bevandic2022multi].
@@ Appendix E Model Structure Details
MSeg (B) [lambert2020mseg] HRNet-W48 (65 M) HRNet-Seg 1024 1024
GroupViT (S) [xu2022groupvit] ViT-S/16 (22 M) GroupViT Transformer 224 224
LSeg+ (B) [li2022lseg] CLIP ViT-B/16 (86 M) DenseCLIP CLIP 512 512
ZegFormer (B) [ding2022zegformer] CLIP ViT-B/16 (86 M) ZegFormer CLIP 640 640
OpenSeg (B) [ghiasi2022openseg] ResNet-101 (45 M) OpenSeg CLIP/ALIGN 640 640
OVSeg (B) [liang2023ovseg] CLIP ViT-B/16 (86 M) MaskFormer CLIP 640 640
MaskCLIP (L) [dong2023maskclip] CLIP ViT-L/14 (307 M) MaskCLIP CLIP 1024 1024
X-Decoder [zou2023xdecoder] DaViT-L (196 M) X-Decoder CLIP 1024 1024
OpenSeeD [zhang2023openseed] Swin-L (197 M) MaskDINO UniCL 1024 1024
SEEM [zou2023seem] DaViT-L (196 M) X-Decoder CLIP 800 1333
APE (B) [shen2024ape] ViT-L (307 M) DETA CLIP 1024 1024
PolyFormer (L) [liu2023polyformer] Swin-L (197 M) PolyFormer BERT 1024 1024
@@
PixelLM [ren2024pixellm] CLIP ViT-L/14 (307 M) PixelLM LLaMA2-13B 448 448
LISA [lai2024lisa] SAM ViT-H (636 M) SAM Vicuna-7B 1024 1024
GLaMM [rasheed2024glamm] SAM ViT-H (636 M) SAM Vicuna-7B 1024 1024
u-LLaVA [xu2024ullava] SAM ViT-H (636 M) SAM Vicuna-7B 1024 1024
EVF-SAM [zhang2024evfsam] SAM ViT-H (636 M) SAM BEiT-3 1024 1024
EVF-SAM2 [zhang2024evfsam] SAM2 Hiera-L (224 M) SAM2 BEiT-3 1024 1024
@@ E.1 Possible Text Encoder Alternatives
VLMo-B/L [bao2022vlmo] joint enc. TFM 768 230/341M same interface as BEiT-3; smaller model; multilingual closest to BEiT-3 $\rightarrow$ just replace tokenizer + dimension in the projector; keep tie-breakers unchanged 18.6/25.4
OFA-B/L [wang2022ofa] joint enc. TFM 768 184/312M instruction-tuned; handy if we ever expose captioning adjust tokenizer and change input dim in projector; reports slightly weaker alignment than BEiT-3 17.9/24.7
Florence-2-Base [xiao2024florence] joint enc. TFM 1024 230M SOTA zero-shot retrieval; 10-lang support none beyond changing tokenizer and input dim in projector 26.3
@@
CLIP-ViT-L/14 [radford2021clip] ViT+Text enc. 768 304M unlimited vocabulary; tiny latency; many open checkpoints semantic vector is not image-conditioned $\rightarrow$ our ablation saw weaker performance. Mitigation: add a 2-layer cross-attn adapter that re-injects image tokens before the projector; expect AP drop if no cross-attn 19.0
EVA-CLIP-E [sun2023eva] ViT-G/14 + Text 1024 610M stronger semantics than CLIP-L memory heavy; expect AP drop if no cross-attn 37.2
SigLIP-2-S [tschannen2025siglip] ViT / Text 512 86M edge-friendly; multilingual expect AP drop if no cross-attn 8.1
@@
CoCa-Base [yu2022coca] ViT enc. + TFM dec. 768 365 M better long-tail semantics need to tap the unimodal decoder hidden state 23.7
PaLI-3B [chen2022pali] ViT-E enc. + T5 dec. 1024 3.0 B 100-lang OCR; robust semantics memory heavy; need to tap the unimodal decoder hidden state 56.4
@@
Kosmos-2 [peng2023kosmos] ViT enc. + LLM dec. 768 1.6B optional box tokens for UX studies Requires a one-step decode per prompt (latency) and an additional MLP to strip location bias 34.8
@@ F.1 Effect of varying tie-breaker tokens
We set the hyperparameter $K=20$, meaning for each prompt (e.g., a category name), our model can identify up to 20 distinct objects. For crowded scenes containing more than 20 objects per category, increasing $K$ is straightforward and advisable. In practice, COCO images typically contain a moderate number of distinct categories and instances (the original COCO paper reports “on average, our dataset contains 3.5 categories and 7.7 instances per image.” [lin2014coco]). The chosen value should match or exceed the maximum expected number of objects per category. For reference, DETR [carion2020detr] used 100 total queries, aligning roughly with the maximum number of objects per image. Our choice ($K=20$) results, on average, in approximately 70 queries per image (20 queries $\times$ 3.5 categories), providing ample coverage for typical scenes.
Further, [zhang2022expected] observed that increasing queries initially improved Average Precision (AP), but then plateaued or even slightly declined when queries became excessive, indicating redundancy in higher query counts. However, recall does improve with more queries, since more detection slots increase the chance to find each object.

### v4

@@ 4.3 Referring Expression Segmentation Performance Analysis
Table 3: Referring segmentation performance (cIoU) comparison on RefCOCOg benchmark validation set between our proposed OpenWorldSAM and prior SOTA methods. We abbreviate the datasets: C (COCO), RC (RefCOCO/+), RCg (RefCOCOg), PL (PACO-LVIS), O365 (Objects365), V (Video segmentation datasets), OID (OpenImages Detection), VG (Visual Genome), ADE (ADE20K), $P$ (PASCAL-Part), PC (PASCAL-VOC). We compare model trainable parameters, model capabilities (OV seg (open-vocabulary segmentation) and Inter Seg (interactive segmentation)), and training data required. “*” denotes an estimate of the trainable parameters, since these models use LoRA [21] with rank-8/16 adapters for finetuning.
@@
X-Decoder (L) [80] CLIP-B [44] (63M) 38 (M) C, RCg, Cap4M 64.6
SEEM (L) [81] CLIP-B [44] (63M) 39 (M) C, RC, RCg, PL 65.6
PolyFormer (L) [36] BERT-B [14] (104M) 450 (M) RC, RCg 71.2
UNINEXT (H) [67] BERT-B [14] (104M) 673 (M) C, RC, O365, V 74.4
APE-B (L) [52] CLIP-L [44] (123M) 42 (M) C, PC, O365, OID, VG, RC, RCg 63.5
PixelLM [48] LLaMA2 [55] (13B) 29 (M) C, RC, ADE, PL, MUSE 69.3
LISA [24] Vicuna [77] (7B) 32 (M) C, RC, ADE, PL, PP 66.4
GLaMM [45] Vicuna [77] (7B) 40 (M) RC, GranD 74.2
u-LLaVA [66] Vicuna [77] (7B) 44 (M) C, RC, ADE, PL, PC 71.6
u-LLaVA [66] Vicuna [77] (7B) 7 (B) C, RC, ADE, PL, PC 74.8
Sa2VA [72] InternVL2 [8] (1B) 22 (M) RC, RCg, V, GranD 72.3
Sa2VA [72] InternVL2 [8] (4B) 35 (M) RC, RCg, V, GranD 74.1
SAMWISE [12] RoBERTa [38] (125M) 5.6 (M) RC 67.3
EVF-SAM [75] BEIT-3-L [59] (0.7B) 674 (M) RC 77.0
OpenWorldSAM BEIT-3-L [59] (0.7B) 4.5 (M) C, RCg 74.0
@@
Table 4: Ablation on the VLM choice, e.g., CLIP [44] model from OpenAI. ‘✔’ denotes modality used during training, and ‘✘’ unused. Only the adapter modules are trainable, and the VLMs are kept frozen. Late fusion means we concatenate text/image features from the last layers of CLIP’s text/image encoder. Early fusion means BEiT-3 processes both modalities in all 24 Transformer layers.
@@ 4.4 Ablation Studies
Multi-modal encoder analysis. In Table 4, we compares performances using different VLM encoders and fusion methods (early fusion vs. late fusion). BEiT-3’s early cross-modal fusion (joint text-image processing across all layers) outperforms CLIP’s late fusion (last-layer concatenation) by +33.9 mIoU, +21.2 PQ, and +13.3 AP on ADE-150, demonstrating that deep semantic integration is critical for aligning language concepts with visual regions, echoing findings by EVF-SAM [75].
@@ A.1 Pre-training
We implement our model in PyTorch, building on the Detectron2 [61] framework. We initialize the base models with the public weights of SAM2-Hiera-Large [^1] https://github.com/facebookresearch/sam2 and BEIT-3-Large [^2] https://huggingface.co/YxZhang/evf-sam2-multitask. The model is pre‑trained for 25 epochs on the COCO‑2017 training split (104K images) [35]. We use the panoptic annotations, which provide pixel‑accurate masks and category labels for all 132 thing and stuff classes. Training is conducted on a single NVIDIA A100 (80 GB) GPU with a batch size of 8. Optimization employs AdamW (learning rate 1e-4). A step decay scheduler drops the learning rate by a factor of 0.1 at 89% and 96% of the total iterations. Compared with recent generalist models, our recipe is markedly more data‑efficient (see Table 6).
@@
X-Decoder [80] COCO, RefC Cap4M 32, 1024 200M (50 Ep $\times$ 4M Img)
OpenSeeD [73] COCO, O365 32, 64 48M (30 Ep $\times$ 1.8M Img)
APE (B) [52] COCO, LVIS, O365, OID, VG, RefC 16 17.28M (16 Bs $\times$ 1.08M Iter)
@@ A.2 Zero-Shot Evaluation
We evaluate semantic, instance, and panoptic segmentation in a zero‑shot setting. For instance segmentation and panoptic segmentation, we apply confidence-score filtering to remove masks with scores below 0.7, followed by non‑maximum suppression (NMS) with IoU threshold 0.5 to remove duplicate detections and retain distinct object instances. The confidence scores, originally termed “estimated IoU scores” in SAM [23, 46], are direct outputs from SAM2’s mask decoder. These scores were optimized during SAM2’s pre-training to select high-quality (i.e., confident) mask outputs.
@@ Appendix C Additional Zero‑Shot Qualitative Results
Figure 8 showcases multiple challenging indoor scenes drawn from ADE20K-150/857 [79], and PASCAL Context-459 [41]. In each sub-panel, we compare example outputs of OpenWorldSAM with comparisons to X-Decoder under both global-matching and oracle-prompts evaluation protocols.
@@
Figure 8: Qualitative comparisons between X-Decoder [80] and OpenWorldSAM on ADE20K-150, ADE20K-857, and PASCAL Context-459.
@@ Appendix D Limitations - Outdoor Generalization
Despite strong results on indoor and everyday photographs, OpenWorldSAM under-performs on driving datasets such as Cityscapes [11] and BDD10K [70] (Table 8). Fine-tuning on Cityscapes narrows the gap, yet performance still trails methods explicitly exposed to multi-domain data. Understanding the source of this shortfall is essential for future extensions.
@@
X-Decoder (L) [80] zero-shot 52.0 24.9 38.1 47.2 17.8
@@
    Domain shift. COCO images are mostly handheld and indoor, whereas Cityscapes/BDD10K contain forward-looking dash-cam frames with motion blur, glare and night scenes. X-Decoder was co-trained on web-scale image-text pairs that include many outdoor photos, so its visual encoder has wider coverage. Large-scale multi-domain training is known to mitigate domain shift [2].
@@ Appendix E Model Structure Details
MSeg (B) [26] HRNet-W48 (65 M) HRNet-Seg 1024 1024
GroupViT (S) [64] ViT-S/16 (22 M) GroupViT Transformer 224 224
LSeg+ (B) [28] CLIP ViT-B/16 (86 M) DenseCLIP CLIP 512 512
ZegFormer (B) [15] CLIP ViT-B/16 (86 M) ZegFormer CLIP 640 640
OpenSeg (B) [19] ResNet-101 (45 M) OpenSeg CLIP/ALIGN 640 640
OVSeg (B) [34] CLIP ViT-B/16 (86 M) MaskFormer CLIP 640 640
MaskCLIP (L) [16] CLIP ViT-L/14 (307 M) MaskCLIP CLIP 1024 1024
X-Decoder [80] DaViT-L (196 M) X-Decoder CLIP 1024 1024
OpenSeeD [73] Swin-L (197 M) MaskDINO UniCL 1024 1024
SEEM [81] DaViT-L (196 M) X-Decoder CLIP 800 1333
APE (B) [52] ViT-L (307 M) DETA CLIP 1024 1024
PolyFormer (L) [36] Swin-L (197 M) PolyFormer BERT 1024 1024
@@
PixelLM [48] CLIP ViT-L/14 (307 M) PixelLM LLaMA2-13B 448 448
LISA [24] SAM ViT-H (636 M) SAM Vicuna-7B 1024 1024
GLaMM [45] SAM ViT-H (636 M) SAM Vicuna-7B 1024 1024
u-LLaVA [66] SAM ViT-H (636 M) SAM Vicuna-7B 1024 1024
EVF-SAM [75] SAM ViT-H (636 M) SAM BEiT-3 1024 1024
EVF-SAM2 [75] SAM2 Hiera-L (224 M) SAM2 BEiT-3 1024 1024
@@ E.1 Possible Text Encoder Alternatives
VLMo-B/L [1] joint enc. TFM 768 230/341M same interface as BEiT-3; smaller model; multilingual closest to BEiT-3 $\rightarrow$ just replace tokenizer + dimension in the projector; keep tie-breakers unchanged 18.6/25.4
OFA-B/L [58] joint enc. TFM 768 184/312M instruction-tuned; handy if we ever expose captioning adjust tokenizer and change input dim in projector; reports slightly weaker alignment than BEiT-3 17.9/24.7
Florence-2-Base [62] joint enc. TFM 1024 230M SOTA zero-shot retrieval; 10-lang support none beyond changing tokenizer and input dim in projector 26.3
@@
CLIP-ViT-L/14 [44] ViT+Text enc. 768 304M unlimited vocabulary; tiny latency; many open checkpoints semantic vector is not image-conditioned $\rightarrow$ our ablation saw weaker performance. Mitigation: add a 2-layer cross-attn adapter that re-injects image tokens before the projector; expect AP drop if no cross-attn 19.0
EVA-CLIP-E [54] ViT-G/14 + Text 1024 610M stronger semantics than CLIP-L memory heavy; expect AP drop if no cross-attn 37.2
SigLIP-2-S [56] ViT / Text 512 86M edge-friendly; multilingual expect AP drop if no cross-attn 8.1
@@
CoCa-Base [71] ViT enc. + TFM dec. 768 365 M better long-tail semantics need to tap the unimodal decoder hidden state 23.7
PaLI-3B [7] ViT-E enc. + T5 dec. 1024 3.0 B 100-lang OCR; robust semantics memory heavy; need to tap the unimodal decoder hidden state 56.4
@@
Kosmos-2 [43] ViT enc. + LLM dec. 768 1.6B optional box tokens for UX studies Requires a one-step decode per prompt (latency) and an additional MLP to strip location bias 34.8
@@ F.1 Effect of varying tie-breaker tokens
We set the hyperparameter $K=20$, meaning for each prompt (e.g., a category name), our model can identify up to 20 distinct objects. For crowded scenes containing more than 20 objects per category, increasing $K$ is straightforward and advisable. In practice, COCO images typically contain a moderate number of distinct categories and instances (the original COCO paper reports “on average, our dataset contains 3.5 categories and 7.7 instances per image.” [35]). The chosen value should match or exceed the maximum expected number of objects per category. For reference, DETR [4] used 100 total queries, aligning roughly with the maximum number of objects per image. Our choice ($K=20$) results, on average, in approximately 70 queries per image (20 queries $\times$ 3.5 categories), providing ample coverage for typical scenes.
Further, [74] observed that increasing queries initially improved Average Precision (AP), but then plateaued or even slightly declined when queries became excessive, indicating redundancy in higher query counts. However, recall does improve with more queries, since more detection slots increase the chance to find each object.